Find the 1st 5 values of the recursive function, then find their 1st and 2nd differences, then state wheather the recursive function has a linear model, quadratic model or neither. f(0)=-2, f(n)=f(n-1)+2n
Answer by drk(1908)

f(0) = -2
f(1) = -2 + 2(1) = 0
f(2) = 0 + 2(2) = 4
f(3) = 4 + 2(3) = 10
f(4) = 10 + 2(4) = 18
f(5) = 18 + 2(5) = 28
-----
first differences: 2, 4, 6, 8, 10
second differences: 2, 2, 2, 2
-----
this is a quadratic model.